How much energy is released during the combustion of 24.0 g of methane, CH4. DeltaH = -90 kJ/mol.


1
Expert's answer
2021-05-26T10:33:03-0400

The combustion of methane undergoes according to the following reaction:

CH4 + 2O2 = CO2 + 2H2O

The number of moles of methane equals:

n(methane) = m(methane) / Mr(methane) = 24.0 g / 16.04 g/mol = 1.5 mol

From here, if 1 mol of methane releases 90 kJ, 1.5 mol of methane releases 1.5 mol × 90 kJ/mol = 135 kJ


Answer: 135 kJ